The area of a rectangle is 100 inches squared. The width is x-15 inches and the length is x inches. Find the dimensions of the rectangle.
Answer by solve_for_x(190)

Area = Length * Width

x(x - 15) = 100

x^2 - 15x - 100 = 0

(x - 20)(x + 5) = 0

x = -5, 20

Since the negative result is impossible, the length is 20 inches.

The corresponding width is x - 15 = 20 - 15 = 5

Solution: Length = 20 inches, Width = 5 inches